OpenStudy (anonymous):

-8A - 3B=0 & 3A - 8B=2 A=6/73 and B=-16/73 how do you solve to find those values for A and B

OpenStudy (anonymous):

The answers for A and B r not given I just know the answer

OpenStudy (anonymous):

-8A-3B=0 -8A=3B A=-3B/8 SUBSTITUTE IN THE OTHER EQUATION 3(-3B/8)-8B=2.........

OpenStudy (anonymous):

Sweet thx

OpenStudy (anonymous):

np :)